# Rambler: Supporting Writing With Speech via LLM-Assisted Macro Revision

### 19 Jan 2024 v1

@@ Abstract.
Dictation enables efficient text input on mobile devices. However, writing with speech can produce disfluent, wordy, and incoherent text and thus requires heavy post-processing. This paper presents Rambler, an LLM-powered graphical user interface that supports gist-level manipulation of dictated text with two main sets of functions: gist extraction and macro revision. Gist extraction generates keywords and summaries as anchors to support the review and interaction with spoken text. LLM-assisted macro revisions allow users to respeak, split, merge and transform dictated text without specifying precise editing locations. Together they pave the way for interactive dictation and revision that help close gaps between spontaneous spoken words and well-structured writing. In a comparative study with 12 participants performing verbal composition tasks, Rambler outperformed the baseline of a speech-to-text editor + ChatGPT, as it better facilitates iterative revisions with enhanced user control over the content while supporting surprisingly diverse user strategies.
copyright: nonejournalyear: 2024conference: CHI ’24: ACM CHI Conference on Human Factors in Computing Systems; May 11-16, 2024; Honolulu, Hawaii
@@ 1. Introduction
Speech, regarded as a natural input modality, has attracted much attention following the breakthroughs of advanced speech recognition (ASR) and natural language processing (NLP). Speech transcription software products like Otter.ai, Google Voice Typing, and Dragon are gaining popularity. Dictation is now an integral keyboard function on most smartphones and tablets. Indeed, dictation can be three to four times faster than typing for text input in the best circumstances (Ruan et al., 2018). This makes it particularly advantageous for inputting long text, such as writing memos or drafting ideas. Mobile devices benefit even more from this given the challenges of using a keyboard and cursor on a touchscreen (Vogel and Baudisch, 2007).
@@
Current interfaces for dictating long text remain largely the same as a traditional text editor with an added microphone. However, writing with speech can be more than just dictation. Previous research (Mehra et al., 2023) showed that speaking differs significantly from typing in terms of both production and memory retention. Transcribed spoken text can be much more verbose, erroneous, and disorganized, leading to high-effort editing (Karat et al., 1999). The verbatim memory of spoken content is also weaker, leading to a likely higher effort in reading the transcripts. Therefore, treating speaking simply as a faster “typing” method can be problematic, especially for long texts. Similarly, researchers for voice user interfaces (VUI) have made long-standing arguments against directly transplanting GUI guidelines to VUI design (Murad et al., 2019). All these call for new affordances and interface solutions for writing with speech. Meanwhile, writing is a highly iterative activity and thought development process. Typically, digital writing requires people’s full attention on a screen while sitting in front of a computer with a mouse and keyboard, but this requires time and an environment, which creates a high barrier to starting. While speech has the potential to be used for faster and lower-barrier writing, its adoption remains limited and the effort to edit resulting transcripts is large.
@@
This work sets out to tackle that and answer the question: Can we identify and build stepping stones towards closing the gap between speaking and writing, by redesigning a dictation interface that caters to the shortcomings of speech? We build on existing solutions in recent works that use NLP and large language models (LLMs) to generate summaries to assist writing (Dang et al., 2022) and review spoken dialogues (Li et al., 2021, 2023a). These interfaces extract the gist of text to support visualization and/or interaction with it, which we call gist-based interfaces. They move beyond automatic text summarization to semantic manipulation, with the interaction centered more around the “gists” of the text rather than the sequences of characters of their words. While such concepts have been explored in traditional writing (Dang et al., 2022), we operationalize it in the context of interacting with spoken text, which is perhaps an even more compelling use case according to the literature showing spoken content remains in memory more as gists than verbatim (Mehra et al., 2023).
@@
We introduce Rambler, an LLM-powered graphical user interface (GUI) that implements a gist-based interface tailored for dictated text (see Figure 1). Users can dictate their spontaneous thoughts and place each segment of recording (i.e., a fragment of an idea) into a “Ramble”. Rambler provides semantic zooming (Perlin and Fox, 1993) to visualize multiple summarization levels of gists for each Ramble, which users can review on demand. Users then focus on iterating on higher-level ideas by manipulating the Rambles, through respeaking, splitting, merging and reorganizing them, which we call macro revision. Behind the scenes, Rambler uses an LLM to automatically clean up the disfluency and punctuation errors from the raw transcript as well as to complete broken sentences and smoothen transitions, which we call micro revision. In addition, Rambler extracts keywords from the text to aid visual skimming across Rambles and semantic zoom levels. Individual keywords can also be activated and deactivated by the user to indicate importance, which then serve as parameters to customize the summary generation and other LLM-based word processing.
@@
Rambler aims to aid the review and revision of spoken text while easing the interaction with text on mobile devices. To evaluate its effectiveness, we conducted a comparative study with 12 participants for a long-form verbal text composition task. The baseline is a standard speech-to-text editor interface with ChatGPT provided on the side. While participants demonstrated varied strategies and feature usage, most of them expressed preference of Rambler over the baseline and acceptance of using it for their own tasks. Their feedback shows Rambler provides better support for developing and reviewing spoken content, supporting iteration while improving user control over content revision.
Our contribution includes the following:
  1. (1)
    Rambler: a novel tool that implements the concept of a gist-based interface to support writing with speech on mobile devices. We motivate our design goals and provide implementation details for Rambler.
  2. (2)
@@
    Empirical findings from a user study revealing the potential of Rambler in helping close the gap between rambling and writing. This study also identified advantages of our specialized LLM-backed GUI over a one-size-fit-all chatbot UI for LLMs, generating insights for designing and building LLM-supported direct manipulation interfaces.
@@ 2.1. Voice Dictation and Editing
While speaking is faster than typing or writing, it can be time-consuming to correct errors caused by the speech artifacts, such as disfluencies and repetitions. Modern speech and NLP research has made much progress in cleaning disfluency and speech recognition errors (Liao et al., 2023; Lou and Johnson, 2020; Tanaka et al., 2018), mostly for post-processing. Researchers have proposed multimodal interaction methods to reduce editing effort, especially on mobile devices. For example, EyeSayCorrect uses gaze to infer users’ editing intentions and allows re-dictation to edit the phrase near eye fixations (Zhao et al., 2022). Similar ideas were applied to other modalities, such as touch (Zhao et al., 2021).
@@ 2.2. Writing with LLMs
While LLMs have made significant progress in the field of writing, there is still room for improvement. LLMs may struggle to retain the writing style and align with the writer’s goal. Ippolito et al. (Ippolito et al., 2022) found that LLMs can struggle to preserve the style and authorial voice of experienced writers. This is supported by Biermann et al.’s (Biermann et al., 2022) study, which found that authors prefer text generation tools that respect their styles and strategies. Rahman et al. (Rahman et al., 2023) studied academic writing with ChatGPT. Their findings revealed that ChatGPT can be useful to generate initial ideas, but falls short in tasks that demand critical thinking, such as literature synthesis, citations, problem statements, research gaps, and data analysis. Taken together, LLMs are a promising tool for writing, but they alone are not yet perfect.
@@ 2.3. Semantic Manipulations on Text
3. Rambler: Gist-based Interface for Writing With Speech
@@ 3.1. Design Rationale and Goals
As explained in the introduction, we aim to design a gist-based dictation interface to support writing with speech by helping users overcome the shortcomings of speech production and memory. The shortcomings identified from previous empirical and theoretical works (Liu et al., 2022; Mehra et al., 2023) can be summarized as follows: 1) speech is spontaneous as it is produced in real-time, so while the gists may be planned, the wording is not; 2) looking at the live transcript while speaking can be distracting, so people tend to have less visual engagement speaking than they would typing, which leads to less micro-revision when speaking; 3) the verbatim memory of spoken content is weaker, which, combined with less visual engagement, may make reviewing content harder.
Figure 2. A labeled screenshot of the Rambler UI. (1) Ramble in default state, with macro-revision interactions accessible through buttons on (2) the top left and (3) the top right. (4) Ramble in re-speaking mode, where voice input is transcribed so that it can be appended to current text, replace the current text, or be discarded using the buttons from left to right at (5). Other macro-revision features are temporarily inaccessible. Fixed at the bottom of the UI is (6), with the semantic merge button, new Ramble button, and semantic zoom slider.
@@ 3.1.1. Rambles as the unit of interaction
With these characteristics, a text editor that transcribes everything linearly will also include all the disfluencies, repetition, and disorganization of raw speech. Although advanced NLP techniques can help correct some of these issues, the user still needs to direct the composition’s meaning and organization. Successful LLM interfaces (e.g.,ChatGPT [^1] https://chat.openai.com/, Claude [^2] https://claude.ai/) have demonstrated the ability to refine unorganized text into high-quality writing. However, such general tools not only rely on users’ ability to prompt the LLM, which was observed to be challenging (Zamfirescu-Pereira et al., 2023b, a), but are also generally hard for users to precisely control where and how their text is edited.
Building on previous works that show the benefits of using NLP or LLM in direct manipulation (Dang et al., 2022; Jiang et al., 2023), we investigate how to leverage recent LLM and NLP techniques to inject intelligent word processing into the GUI for our specific task of writing with speech, with the goal of preserving user control over the their composition. With these considerations, we aim to improve the text editor interface for better interaction with spoken text by establishing the following design goals:
@@
  • Support iterative drafting. Leverage the fast production of speech while recognizing that the first attempt may not be well articulated as written text. Our interface should encourage and support users to easily iterate on drafts by respeaking large chunks of the text before diving into micro-editing.
@@
  • Leverage LLM while preserving user control. The interface should leverage the capacity of LLM in cleaning up, smoothing, and transforming text while allowing users to control where and how the change will be applied.
  • Optimize interactions for mobile devices. Although speech-for-text input can be used on any device, mobile devices particularly benefit from it, given the known challenges of interacting with text on these devices without keyboards. For our evaluation, we target a tablet interface for designing our current interface, as one of the pervasively used mobile devices suitable for long-form writing.
@@ 3.2. Rambler Interface Design
Here, we explain the design of Rambler and its main features supporting macro revision and text reviewing, respectively.
Rambler is built around the creation and manipulation of “Rambles”, illustrated in Figure 1 and 2. Each Ramble is a container (visually separated from other Rambles) that holds its content and multiple buttons for users to interact with. As we envision it, the content of a Ramble roughly corresponds to a single train of thought. Without the pressure of producing a decent piece of text in one go, users may either speak everything into one Ramble and split it at any time, or start a new Ramble when they feel what they will say next is about a different point or subject matter. This should support various writing strategies with or without an outline in mind.
@@ 3.2.1. Supporting text reviewing with clean transcript and gist extraction
Rambler is designed to help users review the dictated text more efficiently with three key features (transcript cleaning, semantic zooming, and keyword spotting) that leverage NLP techniques and LLMs.
Figure 4. Example Rambles before and after (A) semantic merging and (B) semantic splitting. In (A), note that the user must tap to select all the Rambles they wish to include, then execute the semantic merge with the global button (visualized in Figure 2). In (B), the user taps the scissors icon in a Ramble to automatically split it.
Figure 5. Example Ramble showing (a) keyword spotting, with keywords highlighted in light green, and (b) the magic custom prompt modal, with an example prompt.
First, LLMs are used to automatically clean Ramble transcripts when they are added or edited, in order to minimize small speech-to-text (STT) typos (like messy punctuation or misheard words) that impact reading. To avoid distraction, our interface only shows raw speech for real-time feedback when the user is actively dictating; if the user stops dictating, the interface immediately cleans the spoken text, and only displays the clean transcript.
Second, we implement semantic zooming through a slider that controls the degree of summarization at which we view all the Rambles in the composition. As shown in Figure 2.6 and Figure 3, users can move the slider at the bottom of the UI to view text summaries at different levels, specifically at 50%, 25% and 10% of the original length. [^3] Rambler’s prompts for these summary levels were tuned to achieve these levels, but are not explicit to these percentages; see A.2.4. We design this feature to help visualize the core ideas of lengthy text and to allow users to quickly gain a sense of how their ideas should be re-organized and iterated upon.
Last, to aid visual skimming across Rambles, Rambler automatically highlights keywords after a Ramble is added or edited (see Figure 4(a)). Words in the Ramble can be toggled by the user with a tap to be selected/deselected as a keyword. Besides being a visual aid for continuity across zoom levels, users can also tap a regenerate button on a Ramble to re-summarize using their transcript text and their modified keyword selection. This is to address an issue we discovered that summaries generated by LLMs may not be aligned with the users’ idea of Ramble content.
@@ 3.2.2. Supporting macro revision with LLM assistance
Figure 6. Rambler System Architecture Diagram. (A) The user provides speech input to be processed or provides touch input to the interface. (B) Rambler user interface. (C) Our cloud server mediates between user requests and the OpenAI API. (D) The AssemblyAI API is used for transcription and OpenAI GPT-4 API for LLM functionality.
We distinguish between conceptual-level edits, which we call macro-revisions, from fine-grained edits, which we call micro-revisions. Macro-revisions operate on ideas, and in our prototypes happen at a Ramble level – consider functionalities like reorganizing Rambles, and merging or splitting Rambles. Micro-revisions, in contrast, operate on a much smaller scale and, in our prototypes, happen primarily through keyboard editing – consider fixing a single typo or some word choice in a sentence. We consider macro-revisions and micro-revisions complementary, thus our interface also provides keyboard editing, activated by clicking “Edit Text.”
We designed five macro-revision operations: Ramble respeaking, Ramble re-ordering through drag and drop, Ramble merging (concatenation of text) by dragging one Ramble onto another, Ramble splitting (by pressing return during keyboard editing) and Ramble transformation. Although macro-revisions can also be done in a normal text editor, we provide an interface with stronger affordance for doing so, with the goal of encouraging and supporting users to apply these operations on spoken text. We carefully designed a re-speaking mechanism and interface (Figure 2.4), which is activated when the user re-activates the mic of an existing Ramble. The user talks further and then chooses whether to append the new text to the Ramble, replace the old text in the Ramble with it, or discard the new text (Figure 2.4). During the respeaking process, the original text remains above in grey color as a reminder. We do so to aid the user’s memory, as in our design process we learned that without the original text, it is hard to remember what to say.
While all of the macro revisions can be performed manually, we provide an LLM-assisted version for three of the operations – Ramble merge, split, and transformation – and named them semantic merge, semantic split and magic custom prompt.
Semantic merging utilizes the LLM to intelligently merge the content of multiple selected Rambles into the content of one Ramble by clicking a “Semantic Merge” Button (see Figure 2.6). The semantic split button is placed on individual Rambles, which prompts the LLM to divide one particular Ramble into N Rambles based on its content. These two mechanisms are visualized in Figure 4 with example text. The magic custom prompt function is triggered by a magic wand button on a Ramble. It opens up the possibility for expert LLM users to define any custom transformation on Ramble content by directly inputting an LLM prompt (see Figure 4(b)).

### 08 Mar 2024 v2

@@ Abstract.
Dictation enables efficient text input on mobile devices. However, writing with speech can produce disfluent, wordy, and incoherent text and thus requires heavy post-processing. This paper presents Rambler, an LLM-powered graphical user interface that supports gist-level manipulation of dictated text with two main sets of functions: gist extraction and macro revision. Gist extraction generates keywords and summaries as anchors to support the review and interaction with spoken text. LLM-assisted macro revisions allow users to respeak, split, merge, and transform dictated text without specifying precise editing locations. Together they pave the way for interactive dictation and revision that help close gaps between spontaneously spoken words and well-structured writing. In a comparative study with 12 participants performing verbal composition tasks, Rambler outperformed the baseline of a speech-to-text editor + ChatGPT, as it better facilitates iterative revisions with enhanced user control over the content while supporting surprisingly diverse user strategies.
dictation, speech, speech-to-text, STT, text composition, writing, LLM, AI
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642217isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Interaction paradigmsccs: Human-centered computing Interaction techniquesccs: Human-centered computing Natural language interfaces
@@ 1. Introduction
Speech, regarded as a natural input modality, has attracted much attention following the breakthroughs of automatic speech recognition (ASR) and natural language processing (NLP). Speech transcription software products like Otter.ai, Google Voice Typing, and Dragon are gaining popularity. Dictation is now an integral keyboard function on most smartphones and tablets. Indeed, dictation can be three to four times faster than typing for text input under the best circumstances (Ruan et al., 2018). This makes it particularly advantageous for inputting long text, such as writing memos or drafting ideas. Mobile devices benefit even more, given the challenges of using a keyboard and cursor on a touchscreen (Vogel and Baudisch, 2007).
@@
Current interfaces for dictating long text largely mirror traditional text editors with an added microphone. However, writing with speech can be more than just dictation. Previous research (Mehra et al., 2023) showed that speaking differs significantly from typing in terms of both production and memory retention. Transcribed spoken text can be much more verbose, erroneous, and disorganized, leading to high-effort editing (Karat et al., 1999). The verbatim memory of spoken content is also weaker, leading to a likely higher effort in reading the transcripts. Therefore, treating speaking simply as a faster “typing” method can be problematic, especially for long texts. Similarly, researchers of voice user interfaces (VUI) have long argued against directly transplanting GUI guidelines to VUI design (Murad et al., 2019). All these call for new affordances and interface solutions for writing with speech. Further, writing is a highly iterative activity and thought development process. Typically, digital writing requires people’s full attention on a screen while sitting in front of a computer with a mouse and keyboard, but this requires time and an environment, which creates a high barrier to starting. While speech has the potential to be used for faster and lower-barrier writing, its adoption remains limited and the effort to edit resulting transcripts is large.
@@
This work sets out to tackle that challenge and answer the question: Can we identify and build stepping stones towards closing the gap between speaking and writing, by designing a dictation interface that addresses the challenges inherent in speech? We build on existing solutions in recent publications that use NLP and large language models (LLMs) to generate summaries to assist writing (Dang et al., 2022) and review spoken dialogue (Li et al., 2021, 2023a). These interfaces extract the gist of text to support visualization and/or interaction with that text, a paradigm we call gist-based interfaces. They move beyond automatic text summarization to semantic manipulation, with the interaction centered around the “gist” of each piece of text, rather than the sequences of characters in each word of that text. While such concepts have been explored in traditional writing (Dang et al., 2022), we operationalize these ideas in the context of interacting with spoken text, which is perhaps an even more compelling use given the literature that shows spoken content more often remains in memory in the form of gists, rather than verbatim (Mehra et al., 2023).
@@
We introduce Rambler, an LLM-powered graphical user interface (GUI) that implements a gist-based interface tailored to dictated text (see Figure 1). Users can dictate their extemporaneous thoughts and place each segment of recording (i.e., a fragment of an idea) into a Ramble. Rambler provides Semantic Zoom (Perlin and Fox, 1993) to visualize multiple summarization levels in the form of gists for each Ramble, which users can review on demand. Users then focus on iterating on higher-level ideas by manipulating the Rambles, through respeaking, splitting, merging and reorganizing them, activities that we collectively call macro revision. Behind the scenes, Rambler uses an LLM to automatically clean up any disfluencies and punctuation errors from the raw transcript, as well as to complete broken sentences and to smooth transitions, activities that we call micro revision. In addition, Rambler extracts keywords from text to aid visual skimming across Rambles and Semantic Zoom levels. Individual keywords can also be activated and deactivated by the user to indicate importance; these keywords then serve as parameters to customize summary (gist) generation and other LLM-based word processing.
@@
Rambler aims to aid the review and revision of spoken text and ease interactions with text on mobile devices. To evaluate Rambler’s effectiveness, we conducted a comparative study with 12 participants using a long-form verbal text composition task. The baseline is a standard speech-to-text editor interface with ChatGPT provided on the side. While participants demonstrated varied strategies and feature usage, most of them expressed a preference of Rambler over the baseline and acceptance of using it for their own tasks. Their feedback shows Rambler provides better support for developing and reviewing spoken content, supporting iteration, and improving user control over content revision.
@@
Our work makes the following contributions: First, Rambler: a novel tool that implements the concept of a gist-based interface to support writing with speech on mobile devices. We motivate our design goals and provide implementation details for Rambler. Second, we contribute empirical findings from a user study revealing the potential of Rambler in helping close the gap between rambling and writing. Our study also reveals several advantages of our specialized LLM-backed GUI over a one-size-fits-all dialogue-based chatbot UI for LLMs; we thus also offer insights, based on our findings, for designing and building LLM-supported direct manipulation interfaces.
@@ 2.1. Voice Dictation and Editing
Although speaking is faster than typing or writing, it can be time-consuming to correct errors caused by speech artifacts, such as disfluencies and repetitions. Modern speech and NLP research has made much progress in cleaning disfluency and speech recognition errors (Liao et al., 2023; Lou and Johnson, 2020; Tanaka et al., 2018), mostly for post-processing. Researchers have also proposed multimodal interaction methods to reduce editing effort, especially on mobile devices. For example, EyeSayCorrect uses gaze to infer users’ editing intentions and allows re-dictation to edit the phrase near eye fixations (Zhao et al., 2022). Similar ideas were applied to other modalities, such as touch (Zhao et al., 2021).
@@ 2.2. Writing with LLMs
While LLMs have made significant progress in the field of writing, there is still room for improvement. LLMs may struggle to retain the writing style and align with the writer’s goal. Ippolito et al. (Ippolito et al., 2022) found that LLMs can struggle to preserve the style and authorial voice of experienced writers. This is supported by Biermann et al.’s (Biermann et al., 2022a) study, which found that authors prefer text generation tools that respect their styles and strategies. Rahman et al. (Rahman et al., 2023) studied academic writing with ChatGPT. Their findings revealed that ChatGPT can be useful to generate initial ideas, but falls short in tasks that demand critical thinking, such as literature synthesis, citations, problem statements, research gaps, and data analysis. Taken together, LLMs are a promising tool for writing, but they alone are not yet perfect.
@@ 2.3. Semantic Manipulations on Text
3. Rambler: A Gist-based Interface for Writing With Speech
@@ 3.1. Design Rationale and Goals
As explained in the introduction, we aim to design a gist-based dictation interface to support writing with speech, by helping users overcome the challenges of speech production and shortcomings of memory. The challenges and shortcomings identified in previous empirical and theoretical work (Liu et al., 2022; Mehra et al., 2023) can be summarized as follows: 1) speech is spontaneous as it is produced in real-time, so while the gists may be planned, the wording is not; 2) looking at a live transcript while speaking can be distracting, so people tend to have less visual engagement speaking than they would typing, which leads to less micro-revision when speaking; 3) remembering spoken content verbatim is a challenge memory task; combined with reduced visual engagement, these challenges may also make reviewing content harder.
Figure 2. A labeled screenshot of the Rambler UI. (1) Ramble in default state, with revision functions accessible through buttons on (2) and (3). (4) Ramble in re-speaking mode, where voice input is transcribed so that it can be appended to current text, replace the current text, or to be discarded using the buttons on (5). Fixed at the bottom of the UI is (6), with the Semantic Merge button, New Ramble button, and Semantic Zoom slider.
@@ 3.1.1. Rambles as the unit of interaction
With these characteristics, a text editor that transcribes everything linearly will also include all the disfluencies, repetition, and disorganization of raw speech. Although advanced NLP techniques can help correct some of these issues, the user still needs to direct the composition’s meaning and organization. Successful LLM interfaces (e.g., ChatGPT, [^1] https://chat.openai.com/ Claude [^2] https://claude.ai/) have demonstrated the ability to refine unorganized text into high-quality writing. However, such general tools not only rely on users’ ability to prompt the LLM, which has been observed to be challenging (Zamfirescu-Pereira et al., 2023b, a), but also make it harder for users to precisely control where and how their text is edited.
Building on previous work that shows the benefits of using NLP or LLMs in direct manipulation interfaces (Dang et al., 2022; Jiang et al., 2023), we investigate how to leverage recent LLM and NLP techniques to inject intelligent word processing into the GUI for our specific task of writing with speech, with the goal of preserving user control over the composition. With these considerations, we aim to improve the text editor interface for better interaction with spoken text by establishing the following design goals:
@@
  • Support iterative drafting. Leverage the fast production of speech while recognizing that the first attempt may not be well articulated as written text. Our interface should encourage and support users easily iterating on drafts by respeaking large chunks of the text before diving into micro-editing.
@@
  • Leverage LLMs while preserving user control. The interface should leverage the capacity of LLMs in cleaning up, smoothing, and transforming text while allowing users to control where and how changes will be applied.
  • Optimize interactions for mobile devices. Although speech-tor-text input can be used on any device, mobile devices particularly benefit, given the known challenges of interacting with text on these devices without keyboards. For our evaluation, we target a tablet interface for designing our current interface, as it is a commonly used mobile device suitable for long-form writing.
@@
Figure 4. Example Rambles before and after (a) Semantic Merge and (b) Semantic Split. In (a), the user select all the Rambles to include, then press the Semantic Merge button (shown in Figure 2). In (b), the user taps the scissors icon in a Ramble to ask LLM to split it based on content.
@@ 3.2. Rambler Interface Design
Rambler is built around the creation and manipulation of Rambles, illustrated in Figure 1 and 2. Each Ramble is a container (visually separated from other Rambles) that holds its content and multiple buttons for users to interact with. Our intention is for the content of a Ramble to roughly correspond to a single train of thought. Without the pressure of producing a decent piece of text in one shot, users may either speak everything into one Ramble and split it later, or start a new Ramble when they feel what they will say next is about a different point or subject matter. Rambler should support various writing strategies, including those starting both with or without an outline in mind. Here, we explain the design of Rambler and its main features supporting macro revision and text reviewing, respectively.
@@ 3.2.1. Supporting text reviewing with clean transcript and gist extraction
Rambler is designed to help users review dictated text more efficiently with three key features (transcript cleaning, semantic zooming, and highlighting keywords) that leverage NLP and LLMs.
First, Rambler uses an LLM to automatically clean Ramble transcripts when they are added or edited, in order to minimize small speech-to-text (STT) typos (like messy punctuation or misheard words) that impact reading. To avoid distraction, our interface only shows raw speech for real-time feedback when the user is actively dictating; if the user stops dictating, the interface immediately cleans the spoken text, and only displays the clean transcript.
Second, we implement Semantic Zoom using a slider that controls the degree of summarization shown across all the Rambles in a composition. As shown in Figures 2 (component 6) and 3, users can move the slider at the bottom of the UI to view text summaries at different levels, specifically, at 50%, 25%, and 10% of the original length. [^3] Rambler’s prompts for these summary levels were tuned to consistently achieve these levels, but are not explicit about these percentages; see appendix A.2.4. We designed this feature to help visualize the core ideas of lengthy text and to allow users to quickly gain a sense of how their ideas should be re-organized and iterated upon.
Last, to aid visual skimming across Rambles, Rambler automatically highlights keywords after a Ramble is added or edited (see Figure 5-top). Users can also tap any word in a Ramble to manually deselect or select the word as a keyword. Keywords are not only a visual aid for continuity across zoom levels; users can also tap the regenerate button on a Ramble to re-summarize their transcript text based on their modified keyword selection. This helps address an issue we discovered where summaries generated by LLMs may not be aligned with the users’ main ideas in a Ramble.
@@ 3.2.2. Supporting macro revision with LLM assistance
We distinguish between conceptual-level edits, which we call macro-revisions, from fine-grained edits, which we call micro-revisions. Macro-revisions operate on ideas, and in our prototypes happen at a Ramble level – consider functionalities like reorganizing Rambles, and merging or splitting Rambles. Micro-revisions, in contrast, operate on a much smaller scale and, in our prototypes, happen primarily through keyboard editing – consider fixing a single typo or some word choice in a sentence. We consider macro-revisions and micro-revisions complementary, thus our interface also provides keyboard editing, activated by clicking the edit button in a Ramble.
We designed five macro-revision operations: Ramble respeaking, Ramble re-ordering through drag and drop, Ramble merging (concatenation of text) by dragging one Ramble onto another, Ramble splitting (by pressing return during keyboard editing), and Ramble transformation. Although macro-revisions can also be done in a normal text editor, we provide an interface with a stronger affordance for doing so, with the goal of encouraging and supporting users to apply these operations on spoken text. We carefully designed a re-speaking mechanism and interface (Figure 2, zone 4), which is activated when the user re-activates the mic of an existing Ramble. The user talks further and then chooses whether to append the new text to the Ramble, replace the old text in the Ramble with it, or discard the new text (Figure 2, zone 4). During the respeaking process, the original text remains above in grey color as a reminder. We do so to aid the user’s memory, as in our design process we learned that without the original text, it is hard to remember what to say.
While all of the macro revisions can be performed manually, we provide an LLM-assisted version for three of the operations – Ramble merge, split, and transformation – and named them Semantic Merge, Semantic Split, and Magic Custom Prompt. Semantic merging utilizes the LLM to intelligently merge the content of multiple selected Rambles into the content of one Ramble by clicking a Semantic Merge button (see Figure 2.6). Semantic Split, represented by a scissors button, is placed on individual Rambles, which prompts the LLM to divide one particular Ramble into N Rambles based on its content. These two mechanisms are visualized in Figure 4 with example text. The Magic Custom Prompt feature is triggered by a magic wand button on a Ramble. It opens up the possibility for expert LLM users to define any custom transformation on Ramble content by directly inputting an LLM prompt (see Figure 5).
Figure 5. Example transforming a ramble with Magic Custom Prompt: (top) keywords highlighted in light green; (bottom) custom prompt window triggered by the magic wand button on a Ramble. The user inputs an example prompt, and optionally ticks the checkbox for including keywords as context.